# Markov chain Monte Carlo without evaluating the target: an auxiliary variable approach

### v2

@@ 4.2 Comparing the transition density
We use the shorthand notation $d_{\sf TV}(\theta,\theta^{\prime},\omega_{1}):=d_{\sf TV}\left(\mathbb{P}_{\theta,\theta^{\prime}}(\omega_{2}\mid\omega_{1}),\mathbb{P}_{\theta^{\prime},\theta}(\omega_{2}\mid\omega_{1})\right)$. We also define $d_{\sf TV}(\theta,\theta^{\prime})=\sup_{\omega_{1}}d_{\sf TV}(\theta,\theta^{\prime},\omega_{1})$. Moreover, define the largest KL-divergence $\tilde{d}_{\sf KL}(\theta,\theta^{\prime}):=\sup_{\omega_{1}}d_{\sf KL}\left(\mathbb{P}_{\theta,\theta^{\prime}}(\omega_{2}\mid\omega_{1})||\mathbb{P}_{\theta^{\prime},\theta}(\omega_{2}\mid\omega_{1})\right)$, and its symmetrised version $d_{\sf KL}(\theta,\theta^{\prime}):=0.5\sup_{\omega_{1}}\left(d_{\sf KL}\left(\mathbb{P}_{\theta,\theta^{\prime}}(\omega_{2}\mid\omega_{1})||\mathbb{P}_{\theta^{\prime},\theta}(\omega_{2}\mid\omega_{1})\right)+d_{\sf KL}\left(\mathbb{P}_{\theta^{\prime},\theta}(\omega_{2}\mid\omega_{1})||\mathbb{P}_{\theta,\theta^{\prime}}(\omega_{2}\mid\omega_{1})\right)\right).$
@@ 4.3 Spectral gap
The first result is known in Andrieu et al., 2018b . The second result generalises the results of Łatuszyński and Rudolf, (2014). They have been applied in analysing proximal samplers and hybrid slice samplers. See Section 5 of Qin et al., (2023) for details.
@@ 4.5 From algorithms to a more powerful theory, and back
Theorem 2 offers valuable insights for existing algorithms but has significant limitations due to its assumptions. It provides non-trivial lower bounds on the spectral gap only when $\sup_{\theta,\theta^{\prime}}d_{\sf TV}(\theta,\theta^{\prime})<1$. This condition is met by PoissonMH and TunaMH (implicitly), but may not hold in many practical cases. Conversely, the ‘pointwise’ comparison (Theorem 1) is more robust. An appealing research direction is to establish a weaker version of Theorem 2 under relaxed assumptions. Promising approaches include using approximate conductance (Dwivedi et al.,, 2019; Ascolani et al.,, 2024) and weak Poincaré inequalities (Andrieu et al.,, 2022, 2023). These methods have recently shown success in MCMC algorithm analysis. For instance, weak Poincaré inequalities can provide useful bounds on the mixing time for pseudo-marginal MCMC when the spectral gap of the idealised chain is unknown or nonexistent.
A more powerful comparison theory could also lead to new algorithms with theoretical guarantees. As shown in Section 4.4, the assumptions in Zhang and De Sa, (2019) and Zhang et al., (2020) implicitly imply the stringent condition $\sup_{\theta,\theta^{\prime}}d_{\sf TV}(\theta,\theta^{\prime})<1$. Thus, developing a new theory with relaxed assumptions might lead to algorithms with broader applicability.
@@ 5 Numerical experiments
This section presents numerical experiments. We first examine two simulated examples: heterogeneous truncated Gaussian and robust linear regression. We compare our Poisson–Barker and Poisson–MALA algorithms with PoissonMH and full batch algorithms like random-walk Metropolis, MALA, and HMC. In the third experiment, we apply Bayesian logistic regression on the MNIST dataset, comparing TunaMH–SGLD with TunaMH, random-walk Metropolis, MALA, and HMC. Our results show that our gradient-based minibatch methods significantly outperform existing minibatch and full-batch methods.
We implement all methods in Julia. The reproducible code is at https://github.com/ywwes26/MCMC-Auxiliary. In all the figures in this section, the random-walk Metropolis is labeled as MH, while all other methods are labeled by their respective names. Different minibatch algorithms require prior knowledge of different bounds on the model likelihood (refer to the “promise” block in Algorithm 2 and 3). They are derived in supplementary D.
@@ 5.1 Heterogeneous truncated Gaussian
We test PoissonMH, Poisson–MALA, Poisson–Barker, (full batch) random-walk Metropolis, and (full batch) MALA. Following Zhang and De Sa, (2019), the hyperparameter $\lambda$ for the three minibatch algorithms is set to $\lambda=0.0005L^{2}$, resulting in a batch size of about 6000 (6% of the data points). The initial $\theta_{0}$ is drawn from $\mathbb{N}(0,\mathbb{I}_{d})$, and each algorithm’s step size is tuned through several pilot runs to achieve target acceptance rates of 0.25, 0.4, and 0.55. We use Mean Squared Error (MSE) and Effective Sample Size (ESS) to compare these methods. MSEs are calculated for estimating the posterior mean and variance, with true values computed numerically. ESS quantifies the dependence of MCMC samples, with each dimension corresponding to an ESS in an MCMC run. We report the minimum, median, and maximum ESS across all dimensions for all methods.
We compare the clock-time performances of the five methods. Figure 2 shows the MSE over time at different acceptance rates. We find: 1. Poisson–Barker significantly improves PoissonMH, random-walk Metropolis, and MALA for all acceptance rates, both in terms of posterior mean and variance, with more notable improvement at higher acceptance rates. 2. All three minibatch algorithms (PoissonMH, Poisson–MALA, Poisson–Barker) consistently outperform random-walk Metropolis and MALA, demonstrating their effectiveness in the tall data regime. 3. Poisson–MALA matches Poisson–Barker’s best performance at acceptance rates of 0.4 and 0.55 but shows fluctuations and slightly worse performance than PoissonMH at a 0.25 acceptance rate, aligning with Livingstone and Zanella, (2022) that MALA is sensitive to tuning, while Barker’s proposal is more robust.
Table 1 compares ESS per second (ESS/s) across various acceptance rates. Our gradient-based minibatch methods, Poisson–Barker and Poisson–MALA, consistently rank as the top two methods for all acceptance rates (0.25, 0.4, 0.55) and metrics (min, median, max) ESS/s. They improve performance by $1.24-7.02$ times over PoissonMH, $7.61-12.33$ times over MALA, and $13.00-67.24$ times over random-walk Metropolis. At a 0.25 acceptance rate, Poisson-Barker performs best, closely followed by Poisson-MALA. At a 0.55 acceptance rate, Poisson–MALA achieves the highest performance, which is also the highest overall performance when all methods are optimally tuned. At a 0.4 acceptance rate, Poisson–Barker and Poisson–MALA are comparable. Despite the additional gradient estimations, our gradient-based algorithms show substantial benefits from a more efficient proposal.
@@
ESS/s: (Min, Median, Max) Method acceptance rate=0.25 acceptance rate=0.4 acceptance rate=0.55 Best MH (0.04, 0.06, 0.40) (0.03, 0.04, 0.41) (0.03, 0.04, 0.33) (0.04, 0.06, 0.41) MALA (0.06, 0.11, 0.68) (0.09, 0.18, 1.57) (0.09, 0.15, 2.57) (0.09, 0.18, 2.57) PoissonMH (0.37, 0.64, 4.55) (0.31, 0.48, 3.99) (0.25, 0.39, 3.16) (0.37, 0.64, 4.55) Poisson–Barker (0.74, 1.32, 5.77) (0.74, 1.40, 8.38) (0.79, 1.45, 10.68) (0.79, 1.45, 10.68) Poisson–MALA (0.52, 1.09, 5.64) (0.73, 1.37, 12.87) (0.81, 1.53, 22.19) (0.81, 1.53, 22.19)
@@ 5.2 Robust linear regression
We compare Poisson-Barker, Poisson-MALA, PoissonMH, random-walk Metropolis, MALA, and HMC in a robust linear regression example, as considered in Cornish et al., (2019); Zhang et al., (2020); Maclaurin and Adams, (2014). The data is generated following previous works: for $i=1,2,\cdots,N$, covariates $x_{i}\in\mathbb{R}^{d}$ are independently generated from $\mathbb{N}(0,\mathbb{I}_{d})$ and $y_{i}=\sum_{j=1}^{d}x_{ij}+\epsilon_{i}$ with $\epsilon_{i}\sim\mathbb{N}(0,1)$. The likelihood is modeled as $p(y_{i}\mid\theta,x_{i})=\text{Student}(y_{i}-\theta^{\intercal}x_{i}\mid v)$, where $\text{Student}(\cdot\mid v)$ is the density of a Student’s t distribution with $v$ degrees of freedom. The model is “robust” due to the heavier tail of the likelihood function Maclaurin and Adams, (2014). With a flat prior, the tempered posterior is: $\pi(\theta\mid\{y_{i},x_{i}\}_{i=1}^{N})\propto\exp\left\{-\beta\cdot\frac{v+1}{2}\sum_{i=1}^{N}\log\left(1+\frac{(y_{i}-\theta^{\intercal}x_{i})^{2}}{v}\right)\right\}.$
@@
We use $v=4$ as in Cornish et al., (2019); Zhang et al., (2020) and set $d=10$, $N=10^{5}$, $\beta=10^{-4},R=15$. PoissonMH, Poisson–MALA and Poisson–Barker use $\lambda=0.01L^{2}$. A $50$ dimensional example is in supplementary E.2. The number of leapfrog steps in HMC is set to $10$. The initial $\theta_{0}$ is drawn from $\mathbb{N}(0,\mathbb{I}_{d})$, and the step size of each algorithm is tuned to achieve acceptance rates of $0.25,0.4,$ and $0.55$. Our true parameter is $\theta^{\star}=(1,1,\ldots,1)$.
Figure 3 shows the MSE of estimating the true coefficients over time. Poisson–MALA and Poisson–Barker are much more efficient than PoissonMH, and all minibatch methods outperform full-batch methods. Poisson–MALA and Poisson–Barker adapt to the posterior faster due to gradient information. For example, in Figure 3 (a), Poisson–MALA and Poisson–Barker have an MSE much less than 0.25 at 0.05 seconds, whereas PoissonMH is around 0.5. Full-batch methods adapt slowly because they require a complete dataset scan. MALA outperforms random-walk Metropolis when properly tuned. HMC performs worse than MALA and sometimes even worse than random-walk Metropolis due to the high cost of its $10$ leapfrog steps, making it about $20$ times more costly.
We also present the ESS/s comparisons in Table 2. The results have a similar trend to Section 5.1. Poisson–MALA and Poisson–Barker are the top two methods across all metrics, showing improvements ranging from 2.61 to 4.62 times over PoissonMH, and achieving nearly or more than 100 times the performance of full-batch methods.
Supplementary E.1 presents additional experiments that compare different HMC leapfrog steps and MSE over the iteration count. Supplementary E.2 considers a 50-dimensional robust linear regression with other settings unchanged, yielding similar conclusions.
@@
ESS/s: (Min, Median, Max) Method acceptance rate=0.25 acceptance rate=0.4 acceptance rate=0.55 Best MH (1.3, 1.6, 1.8) (1.3, 1.5, 1.7) (0.9, 1.0, 1.1) (1.3, 1.6, 1.8) MALA (1.6, 1.7, 1.8) (3.3, 3.6, 3.8) (3.8, 4.1, 4.3) (3.8, 4.1, 4.3) HMC-10 (0.6, 0.7, 0.7) (0.7, 0.8, 0.8) (0.7, 0.8, 0.8) (0.7, 0.8, 0.8) PoissonMH (73.7, 76.7, 79.5) (66.2, 68.6, 70.7) (41.4, 44.4, 46.3) (73.7, 76.7, 79.5) PoissonMH-Barker (135.5, 142.0, 145.0) (195.1, 199.1, 206.9) (197.4, 204.3, 207.7) (197.4, 204.3, 207.7) PoissonMH-MALA (156.8, 165.4, 170.6) (293.2, 304.3, 312.5) (340.2, 353.0, 361.2) (340.2, 353.0, 361.2)
@@ 5.3 Bayesian logistic regression
We compare TunaMH–SGLD, TunaMH, MALA, HMC, random-walk Metropolis to a Bayesian logistic regression task using the MNIST handwritten digits dataset. We focus on classifying handwritten 3s and 5s, which can often appear visually similar. The training and test set has 11,552 and 1,902 samples, respectively. Figure 5 shows examples of the original data. Following Welling and Teh, (2011); Zhang et al., (2020), we use the first 50 principal components of the $28\times 28$ features as covariates.
@@
For TunaMH and TunaMH–SGLD, we follow Zhang et al., (2020) and set the hyperparameter $\chi=10^{-5}$. The batch size $\lvert B\rvert$ in TunaMH–SGLD is set as $20$, so the computation overhead is minimal. We also clip the gradient norm to 2. The original study tests step sizes of $0.0008$ for TunaMH and $0.004$ for random-walk Metropolis. Therefore, we choose step sizes $\{0.0008,0.001,0.002,0.004\}$ for all algorithms. HMC uses 5 leapfrog steps. More extensive comparisons are provided in the supplementary.
The test accuracy comparisons are shown in Figure 5. Due to an extra minibatch for gradient estimation, TunaMH–SGLD converges faster than TunaMH for all step sizes, especially larger ones. Both of them significantly outperform full batch methods, consistent with our previous observations. HMC converges poorly with the stepsizes considered.
Supplementary E.3 includes more extensive experiments, testing all algorithms with larger step sizes and comparing their performance when optimally tuned. Supplementary E.4 compares these algorithms using Bayesian logistic regression for classifying 7s and 9s.

### v3

@@ 4.2 Comparing the transition density
We use the shorthand notation $d_{\sf TV}(\theta,\theta^{\prime},\omega_{1}):=d_{\sf TV}\left(\mathbb{P}_{\theta,\theta^{\prime}}(\omega_{2}\mid\omega_{1}),\mathbb{P}_{\theta^{\prime},\theta}(\omega_{2}\mid\omega_{1})\right)$. We also define $d_{\sf TV}(\theta,\theta^{\prime})=\sup_{\omega_{1}}d_{\sf TV}(\theta,\theta^{\prime},\omega_{1})$. Moreover, define the largest KL-divergence $\tilde{d}_{\sf KL}(\theta,\theta^{\prime}):=\sup_{\omega_{1}}d_{\sf KL}\left(\mathbb{P}_{\theta,\theta^{\prime}}(\omega_{2}\mid\omega_{1})||\mathbb{P}_{\theta^{\prime},\theta}(\omega_{2}\mid\omega_{1})\right)$, and its symmetrized version $d_{\sf KL}(\theta,\theta^{\prime}):=0.5\sup_{\omega_{1}}\left(d_{\sf KL}\left(\mathbb{P}_{\theta,\theta^{\prime}}(\omega_{2}\mid\omega_{1})||\mathbb{P}_{\theta^{\prime},\theta}(\omega_{2}\mid\omega_{1})\right)+d_{\sf KL}\left(\mathbb{P}_{\theta^{\prime},\theta}(\omega_{2}\mid\omega_{1})||\mathbb{P}_{\theta,\theta^{\prime}}(\omega_{2}\mid\omega_{1})\right)\right).$
@@ 4.3 Spectral gap
The first result is known in Andrieu et al., 2018b . The second result generalizes the results of Łatuszyński and Rudolf, (2014). They have been applied in analyzing proximal samplers and hybrid slice samplers. See Section 5 of Qin et al., (2023) for details.
@@ 4.5 From algorithms to a more powerful theory, and back
Theorem 2 offers valuable insights for existing algorithms but has significant limitations due to its assumptions. It provides non-trivial lower bounds on the spectral gap only when $\sup_{\theta,\theta^{\prime}}d_{\sf TV}(\theta,\theta^{\prime})<1$. This condition is met by PoissonMH and TunaMH (implicitly), but may not hold in many practical cases. Conversely, the ‘pointwise’ comparison (Theorem 1) is more robust. An appealing research direction is to establish a weaker version of Theorem 2 under relaxed assumptions. Promising approaches include using approximate conductance (Dwivedi et al.,, 2019; Ascolani et al.,, 2024) and weak Poincaré inequalities (Andrieu et al.,, 2022, 2023). These methods have recently shown success in MCMC algorithm analysis. For instance, weak Poincaré inequalities can provide useful bounds on the mixing time for pseudo-marginal MCMC when the spectral gap of the idealized chain is unknown or nonexistent.
A more powerful comparison theory could also lead to new algorithms. As shown in Section 4.4, the assumptions in Zhang and De Sa, (2019) and Zhang et al., (2020) implicitly imply the stringent condition $\sup_{\theta,\theta^{\prime}}d_{\sf TV}(\theta,\theta^{\prime})<1$. Thus, developing a new theory with relaxed assumptions might lead to algorithms with broader applicability.
@@ 5 Numerical experiments
We first examine two simulated examples: heterogeneous truncated Gaussian and robust linear regression. We compare our Poisson–Barker and Poisson–MALA algorithms with PoissonMH and full batch algorithms like random-walk Metropolis, MALA, HMC, and Barker Livingstone and Zanella, (2022). In the third experiment, we apply Bayesian logistic regression on the MNIST dataset, comparing Tuna–SGLD with TunaMH, random-walk Metropolis, MALA, HMC, and Barker. Our results show that our gradient-based minibatch methods significantly outperform existing minibatch and full-batch methods.
We implement all methods in Julia, with code at https://github.com/ywwes26/MCMC-Auxiliary. In all the figures, the random-walk Metropolis is labeled as MH, while all other methods are labeled by their respective names. Minibatch algorithms require computing several parameters described in Algorithm 2 and 3 before implementation. They are derived in supplementary E.
@@ 5.1 Heterogeneous truncated Gaussian
We test PoissonMH, Poisson–MALA, Poisson–Barker, (full batch) random-walk Metropolis, (full batch) MALA and (full batch) Barker. Following Zhang and De Sa, (2019), the hyperparameter $\lambda$ for the three minibatch algorithms is set to $\lambda=0.0005L^{2}$, resulting in a batch size of about 6000 (6% of the data points). The initial $\theta_{0}$ is drawn from $\mathbb{N}(0,\mathbb{I}_{d})$, and each algorithm’s step size is tuned through several pilot runs to achieve target acceptance rates of 0.25, 0.4, and 0.55. We use Mean Squared Error (MSE) and Effective Sample Size (ESS) to compare these methods. MSEs are calculated for estimating the posterior mean and variance, with true values computed numerically. ESS quantifies the dependence of MCMC samples, with each dimension corresponding to an ESS in an MCMC run. We report the minimum, median, and maximum ESS across all dimensions for all methods.
We compare the clock-time performances of the five methods. Figure 2 shows the MSE over time at different acceptance rates. We find: 1. Poisson–Barker significantly improves PoissonMH, random-walk Metropolis, MALA and Barker for all acceptance rates, both in terms of posterior mean and variance, with more notable improvement at higher acceptance rates. 2. All three minibatch algorithms (PoissonMH, Poisson–MALA, Poisson–Barker) consistently outperform random-walk Metropolis, MALA and Barker. 3. Poisson–MALA matches Poisson–Barker’s best performance at acceptance rates of 0.4 and 0.55 but shows worse performance than PoissonMH at a 0.25 acceptance rate, aligning with Livingstone and Zanella, (2022) that MALA is sensitive to tuning, while Barker is more robust.
Table 1 compares ESS per second (ESS/s) across various acceptance rates. Our gradient-based minibatch methods, Poisson–Barker and Poisson–MALA, consistently rank as the top two methods for all acceptance rates (0.25, 0.4, 0.55) and metrics (min, median, max) ESS/s. They improve performance by $1.35-7.84$ times over PoissonMH, $4.11-12.17$ times over MALA, $5.90-16.40$ times over Barker, and $15.03-76.53$ times over random-walk Metropolis. At 0.25 and 0.4 acceptance rates, Poisson–Barker and Poisson–MALA are comparable. At a 0.55 acceptance rate, Poisson–MALA achieves the highest performance, which is also the highest overall performance when all methods are optimally tuned. Despite the additional gradient estimations, our gradient-based algorithms show substantial benefits from a more efficient proposal.
@@
ESS/s: (Min, Median, Max) Method acceptance rate=0.25 acceptance rate=0.4 acceptance rate=0.55 Best MH (0.03, 0.06, 0.38) (0.03, 0.05, 0.36) (0.03, 0.04, 0.30) (0.03, 0.06, 0.38) MALA (0.06, 0.12, 0.66) (0.08, 0.17, 1.73) (0.08, 0.18, 2.72) (0.08, 0.18, 2.72) Barker (0.10, 0.17, 0.74) (0.10, 0.20, 1.15) (0.10, 0.19, 1.40) (0.10, 0.20, 1.40) PoissonMH (0.33, 0.58, 4.03) (0.31, 0.50, 4.23) (0.24, 0.36, 2.93) (0.33, 0.58, 4.23) Poisson–Barker (0.73, 1.33, 5.71) (0.81, 1.56, 9.00) (0.81, 1.59, 11.18) (0.81, 1.59, 11.18) Poisson–MALA (0.59, 1.08, 6.25) (0.75, 1.41, 14.21) (0.85, 1.61, 22.96) (0.85, 1.61, 22.96)
@@ 5.2 Robust linear regression
We compare Poisson-Barker, Poisson-MALA, PoissonMH, random-walk Metropolis, MALA, Barker and HMC in a robust linear regression example, as considered in Cornish et al., (2019); Zhang et al., (2020); Maclaurin and Adams, (2014). The data is generated following previous works: for $i=1,2,\cdots,N$, covariates $x_{i}\in\mathbb{R}^{d}$ are independently generated from $\mathbb{N}(0,\mathbb{I}_{d})$ and $y_{i}=\sum_{j=1}^{d}x_{ij}+\epsilon_{i}$ with $\epsilon_{i}\sim\mathbb{N}(0,1)$. The likelihood is modeled as $p(y_{i}\mid\theta,x_{i})=\text{Student}(y_{i}-\theta^{\intercal}x_{i}\mid v)$, where $\text{Student}(\cdot\mid v)$ is the density of a Student’s t distribution with $v$ degrees of freedom. The model is “robust” due to the heavier tail of the likelihood function Maclaurin and Adams, (2014). With a flat prior, the tempered posterior is: $\pi(\theta\mid\{y_{i},x_{i}\}_{i=1}^{N})\propto\exp\left\{-\beta\cdot\frac{v+1}{2}\sum_{i=1}^{N}\log\left(1+\frac{(y_{i}-\theta^{\intercal}x_{i})^{2}}{v}\right)\right\}.$
@@
We use $v=4$ as in Cornish et al., (2019); Zhang et al., (2020) and set $d=10$, $N=10^{5}$, $\beta=10^{-4},R=15$. PoissonMH, Poisson–MALA and Poisson–Barker use $\lambda=0.01L^{2}$. A $50$ dimensional example is in supplementary F.2. The number of leapfrog steps in HMC is set to $10$. The initial $\theta_{0}$ is drawn from $\mathbb{N}(0,\mathbb{I}_{d})$, and the step size of each algorithm is tuned to achieve acceptance rates of $0.25,0.4,$ and $0.55$. Our true parameter is $\theta^{\star}=(1,1,\ldots,1)$.
Figure 3 shows the MSE of estimating the true coefficients over time. Poisson–MALA and Poisson–Barker are much more efficient than PoissonMH, and all minibatch methods outperform full-batch methods. HMC performs worse than MALA and Barker, sometimes even worse than random-walk Metropolis due to the high cost of its $10$ leapfrog steps, making it about $20$ times more costly.
We also present the ESS/s comparisons in Table 2. The results have a similar trend to Section 5.1. Poisson–MALA and Poisson–Barker are the top two methods across all metrics, showing improvements ranging from 1.91 to 8.10 times over PoissonMH, and achieving nearly or more than 100 times the performance of full-batch methods.
Supplementary F.1 presents additional experiments that compare different HMC leapfrog steps and MSE over the iteration count. Supplementary F.2 considers a 50-dimensional experiment with other settings unchanged. The findings are consistent with this experiment.
@@
ESS/s: (Min, Median, Max) Method acceptance rate=0.25 acceptance rate=0.4 acceptance rate=0.55 Best MH (2.3, 2.6, 2.9) (1.6, 1.9, 2.1) (1.3, 1.5, 1.8) (2.3, 2.6, 2.9) MALA (2.4, 2.8, 2.9) (4.2, 4.6, 5.0) (6.0, 6.2, 6.8) (6.0, 6.2, 6.8) Barker (2.3, 2.6, 2.8) (2.6, 2.8, 3.1) (3.4, 3.5, 3.8) (3.4, 3.5, 3.8) HMC-10 (1.0, 1.1, 1.2) (0.9, 1.1, 1.1) (1.1, 1.1, 1.2) (1.1, 1.1, 1.2) PoissonMH (126.4, 133.5, 138.7) (88.4, 95.5, 98.2) (70.7, 74.2, 77.0) (126.4, 133.5, 138.7) PoissonMH-Barker (241.6, 255.7, 264.5) (260.3, 272.9, 281.9) (321.9, 338.8, 348.8) (321.9, 338.8, 348.8) PoissonMH-MALA (285.9, 294.7, 314.0) (415.2, 422.1, 446.4) (572.7, 591.6, 613.4) (572.7, 591.6, 613.4)
@@ 5.3 Bayesian logistic regression
We compare Tuna–SGLD, TunaMH, MALA, Barker, HMC, random-walk Metropolis to a Bayesian logistic regression task using the MNIST handwritten digits dataset. We focus on classifying handwritten 3s and 5s, which can often appear visually similar. The training and test set has 11,552 and 1,902 samples, respectively. Figure 5 shows examples of the original data. Following Welling and Teh, (2011); Zhang et al., (2020), we use the first 50 principal components of the $28\times 28$ features as covariates.
@@
For TunaMH and Tuna–SGLD, we follow Zhang et al., (2020) and set the hyperparameter $\chi=10^{-5}$. The batch size $\lvert B\rvert$ in Tuna–SGLD is set as $20$, so the computation overhead is minimal. We also clip the gradient norm to 2. The original study tests step sizes of $0.0008$ for TunaMH and $0.004$ for random-walk Metropolis. To make the experiment more comprehensive, we choose step sizes ${0.0008,0.001,0.002,0.004}$ for all algorithms. HMC uses 5 leapfrog steps. More extensive comparisons are provided in the supplementary.
The test accuracy comparisons are shown in Figure 5. Due to an extra minibatch for gradient estimation, Tuna–SGLD converges faster than TunaMH for all step sizes, especially larger ones. Both of them significantly outperform full batch methods, consistent with our previous observations. HMC converges poorly with the stepsizes considered.
Supplementary F.3 includes more extensive experiments, testing all algorithms with larger step sizes and comparing their performance when optimally tuned. Supplementary F.4 compares these algorithms using Bayesian logistic regression for classifying 7s and 9s.